OpenStudy (cathyangs):

Is there a useful trick to remembering squares and cubes? Is there a useful trick to remembering squares and cubes? @Mathematics

OpenStudy (cathyangs):

up until 20.

hero (hero):

1,4,9,16,25,36... 1,8,27,64... Write them up to 20 and look at them everyday

OpenStudy (anonymous):

For some numbers yes. for example: 57^2. 7 squared is 49, and 25+7=32, so the answer is 3249. 53^2. 3 squared is 9, 25+3 = 28. so the answer is 2809. this works for numbers in the 50's

hero (hero):

Make a chart

OpenStudy (anonymous):

theres trick for numbers that end in 5. 115^2. 5^2 is 25. 11*12=132. so the answer is 13225. 195^2. 5^2 is 25. 19*20 is 380. so the answer is 38025.

OpenStudy (cathyangs):

thanks guys/girls!

OpenStudy (anonymous):

Numbers close to 100. 103^2. 3^2=9 100+2*3=106. answer is 10609 97^2. 97 is 3 away from 100. 3 squared is 9. 100-2(3)=94. answer is 9409.

hero (hero):

Joe, you should change your name to Math Nerd

OpenStudy (anonymous):

Making a chart might be useful.

OpenStudy (anonymous):

i dont think we can change screen names on here >.< without losing medal stuff anywhos.

OpenStudy (anonymous):

Actually there's no need to use a trick or anything, if you just practice, you'll get them in no time.

OpenStudy (asnaseer):

@cathyangs - if you are really interested in learning this, you may find these useful: 1. http://www.math.hmc.edu/~benjamin/papers/Cubing.pdf 2. http://www.quickermaths.com/shortcut-to-find-the-cube-of-a-number/ 3. http://patternizer.wordpress.com/numbers/math-tricks/

OpenStudy (anonymous):

What interesting is why these tricks work. Its not really a trick, but thinking of the number differently. for example:\[103^2=(100+3)^2=10000+2(3)(100)+3^2\] thats why it works. same with the 50's and numbers that end in 5, etc.

hero (hero):

What about 103^2 = (99 + 4)^2 Does that work?

OpenStudy (anonymous):

it does, but the idea is to use numbers that are easier to work with. Like 100's, 50's etc. But yes that would work :)

hero (hero):

I figured it would

OpenStudy (anonymous):

you could take it as 100 + 3

hero (hero):

Yes, as Joe already mentioned